Here is the question: "For the given event, state whether the difference between what occurred and what you would have expected by chance is statistically significant or not statistically significant. After 36 rolls of two six-sided dice, the numbers on the dice add to 4 two times."
Answer #1

probability of dice add to 4 on a single roll of two dice =3/36=1/12 (since (1,3),(2,2),(3,1) are 3 outcomes where sum is 4)

P(getting 2 or less number of times the sum of 4) =P(X<=2) =P(X=0)+P(X=1)+P(X=2)

=(36C0)*(1/12)0(11/12)36+(36C1)*(1/12)1(11/12)35+(36C2)*(1/12)2(11/12)34=0.4134

since probability of above event or more extreme is greater than 0.05 , the event is not statistically significant.